3. (34 points total) Below are EI-MS and IR data for an unknown compound: 134 100 106 51 5 Mt 31 39 91 118 150 10 20 30 40 50 60 70 80 90 100 110 120 130 140 150 160 170 180 a 3 points) Draw the structure that best fits the data in the box provided.

Help!!!
Answer #1

The structure that best fits the given IR and EI-MS data can be drawn as follows.

IR spectrum:

(i) A signal at ~ 3400cm-1 corresponds to N-H stretch.

(ii) A signal at ~ 1600 cm-1 corresponds to aromatic C=C stretch.

(iii) A signal at ~ 1500 cm-1 corresponds to =C-N stretch.

(iv) A signal at ~ 1050 cm-1 corresponds to a C-O stretch of ether.

Mass spectrum:

(i) For the molecular ion (M+) peak, m/z = 165

(ii) The methyl group fragmentation results in m/z = 150 (i.e. 165-15)

Chemical Formula: CgH12NO Exact Mass 150 Chemical Formula: CH Exact Mass: 15

(iii) The OMe fragmentation results in m/z = 134 (i.e. 165 - 31)

Chemical Formula: CH20 Exact Mass: 31 Chemical Formula: CoH12N Exact Mass 134

(iv) The removal of ethylene from the fragment (m/z = 134) results m/z = 106 (i.e. 134-28)

H2C-CH2 Chemical Formula:CoH12N Exact Mass: 134 Chemical Fomula: C/N Chemical Formula C2H Exact Mass: 106 Exact Mass: 28

(v) The removal of NH=CH2 from the fragment (m/z = 106) results m/z = 77 (106-29)

HN-CH2 Chemical Formula: C7H3N Exact Mass: 106 Chemical Formula: CaH Exact Mass: 77 Chemical Formula: CH N Exact Mass: 29